# Sequence Reconstruction for Single-Deletion Single-Substitution Channel

### v1

@@ IV Proof of Theorem 1
1) If $|T^{X}\cap[i_{1}+1,i_{d}]|=0$, then for each $(\bm{z},\bm{z}^{\prime})\in\Lambda_{1}$, either $\bm{z}=x_{[n]\backslash i_{1}}$ or $\bm{z}^{\prime}=x^{\prime}_{[n]\backslash i_{d}}$. Hence, by Remark 6, we have $\Omega^{L}_{1}=\bigcup_{(\bm{z},\bm{z}^{\prime})\in\Lambda^{L}_{1}}B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})\subseteq\Omega^{L}_{0}.$
2) If $|T^{X}\cap[i_{1}+1,i_{d}]|\neq 0$ and $d=2$, we have $|\Lambda^{L}_{1,1}|=|\Lambda^{L}_{1,3}|=1$ and $|\Lambda^{L}_{1,2}|\leq 1$. Hence, we have $|\Lambda^{L}_{1}|\leq 3$.
3) If $|T^{X}\cap[i_{1}+1,i_{d}]|\neq 0$ and $d\geq 3$, we have $T^{L}\cap[i_{2}+1,i_{d}]\neq\emptyset$ or $T^{L}\cap[i_{1}+1,i_{d-1}]\neq\emptyset$. Then by Claims 1.1 and 1.3, $|\Lambda^{L}_{1,1}|+|\Lambda^{L}_{1,3}|\leq 1$. Moreover, by Claim 1.2, $|\Lambda^{L}_{1,2}|\leq 1$. Thus, we have $|\Lambda^{L}_{1}|=|\bigcup_{i=1}^{3}\Lambda^{L}_{1,i}|\leq 2$.
For $d=2$, to simplify the expressions, we need the following notations. For $X\in\{L,R\}$, let
@@
Now, we prove the upper bound in Theorem 1, that is, if $d=d_{\text{H}}(\bm{x},\bm{x}^{\prime})\geq 2$ and $n\geq\max\{\frac{q+23}{2},\frac{5q+19}{q-1}\}$, then
@@
Suppose $d\geq 3$.
@@
To prove the tightness of this bound, we give two examples of $\bm{x},\bm{x}^{\prime}\in\Sigma_{q}^{n}$ such that $d_{\text{H}}(\bm{x},\bm{x}^{\prime})=2$ and $|B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})|=2qn-3q-2-\delta_{q,2}$.
@@
It is not hard to verify that
@@
By the above discussions, we prove Theorem 1.

### v2

@@ IV Proof of Theorem 1
1) If $|T^{L}\cap[i_{1}+1,i_{d}]|=0$, then for each $(\bm{z},\bm{z}^{\prime})\in\Lambda_{1}$, either $\bm{z}=x_{[n]\backslash i_{1}}$ or $\bm{z}^{\prime}=x^{\prime}_{[n]\backslash i_{d}}$. Hence, by Remark 6, we have $\Omega^{L}_{1}=\bigcup_{(\bm{z},\bm{z}^{\prime})\in\Lambda^{L}_{1}}B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})\subseteq\Omega^{L}_{0}.$
2) If $|T^{L}\cap[i_{1}+1,i_{d}]|\neq 0$ and $d=2$, we have $|\Lambda^{L}_{1,1}|=|\Lambda^{L}_{1,3}|=1$ and $|\Lambda^{L}_{1,2}|\leq 1$. Hence, we have $|\Lambda^{L}_{1}|\leq 3$.
3) If $|T^{L}\cap[i_{1}+1,i_{d}]|\neq 0$ and $d\geq 3$, we have $T^{L}\cap[i_{2}+1,i_{d}]\neq\emptyset$ or $T^{L}\cap[i_{1}+1,i_{d-1}]\neq\emptyset$. Then by Claims 1.1 and 1.3, $|\Lambda^{L}_{1,1}|+|\Lambda^{L}_{1,3}|\leq 1$. Moreover, by Claim 1.2, $|\Lambda^{L}_{1,2}|\leq 1$. Thus, we have $|\Lambda^{L}_{1}|=|\bigcup_{i=1}^{3}\Lambda^{L}_{1,i}|\leq 2$.
For $d=2$, to simplify the expressions, we introduce the following notations. For $X\in\{L,R\}$, let
@@
Now, we can prove Theorem 1.
proof 15 (Proof of Theorem 1)
We first prove that if $d=d_{\text{H}}(\bm{x},\bm{x}^{\prime})\geq 2$ and $n\geq\max\{\frac{q+23}{2},\frac{5q+19}{q-1}\}$, then
@@
We divide our discussions into the following two cases.
Case 1: $d\geq 3$.
@@
To prove the tightness of this bound, we consider the following two examples.
@@
For this example, it is not hard to verify that
@@
By the above discussions, we proved Theorem 1.
@@ V Conclusions and future work
Another interesting problem is to generalize the method to single-deletion $s$-substitution channel, that is, to derive a tight upper bound of $\big{|}B^{\text{D,S}}_{1,s}(\bm{x},\bm{x}^{\prime})\big{|}$, where $s\geq 2$ is any fixed integer. We need to consider the set $\big{\{}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):j,j^{\prime}\in[n]~{}\text{and}~{}d_{\text{H}}(x_{[n]\backslash j^{\prime}},x^{\prime}_{[n]\backslash j})\leq 2s\big{\}}$ and can divide it by the similar method of this paper. Correspondingly, $B^{\text{D,S}}_{1,s}(\bm{x},\bm{x}^{\prime})$ can be divided into some subsets and each subset can be easily determined. However, the difficulty is how to find the intersection of these subsets.